### v1

# Not-Just-Scaling Laws: Towards a Better Understanding of the Downstream Impact of Language Model Design Decisions

Abstract
Improvements in language model capabilities are often attributed to increasing model size or training data, but in some cases smaller models trained on curated data or with different architectural decisions can outperform larger ones trained on more tokens. What accounts for this? To quantify the impact of these design choices, we meta-analyze 92 open-source pretrained models across a wide array of scales, including state-of-the-art open-weights models as well as less performant models and those with less conventional design decisions. We find that by incorporating features besides model size and number of training tokens, we can achieve a relative 3-28% increase in ability to predict downstream performance compared with using scale alone. Analysis of model design decisions reveal insights into data composition, such as the trade-off between language and code tasks at 15-25% code, as well as the better performance of some architectural decisions such as choosing rotary over learned embeddings. Broadly, our framework lays a foundation for more systematic investigation of how model development choices shape final capabilities. [^1] Code and data are available at https://github.com/nightingal3/llm-pretraining-behaviours for the community to build upon.
[…]
Not-Just-Scaling Laws: Towards a Better Understanding of the Downstream Impact of Language Model Design Decisions
Emmy Liu1, Amanda Bertsch1, Lintang Sutawika1, Lindia Tjuatja1, Patrick Fernandes1,2,3 Lara Marinov1, Michael Chen1, Shreya Singhal1, Carolin Lawrence4, Aditi Raghunathan1, Kiril Gashteovski4, 5, Graham Neubig1 1 Carnegie Mellon University, Language Technologies Institute 2Instituto Superior Técnico (Lisbon ELLIS Unit), 3 Instituto de Telecomunicações 4 NEC Laboratories Europe, Germany 5 Center for Advanced Interdisciplinary Research, Ss. Cyril and Methodius Uni. of Skopje, Germany Correspondence: emmy@cmu.edu
Figure 1: We document design decisions from open-weights models related to both architecture and data composition, and train predictors for downstream task performance. This allows us to examine the impact of model design choices individually.
1 Introduction
The effectiveness of language model training depends critically on decisions made during pretraining. For instance, the effectiveness of scaling up data depends on its composition – processing even a trillion tokens would be ineffective if they all consisted of the word “the”. Language model performance has been found to be fairly predictable through scaling laws (Kaplan et al. (2020), section 2) – extrapolations of model performance based on the parameter counts and number of tokens the models were trained on. However, scaling laws based on only these two aspects do not always explain downstream task performance Diaz and Madaio (2024); Isik et al. (2024).
The research community has made progress in understanding how training decisions impact downstream performance with respect to data composition. For instance, controlled studies have demonstrated that training on code data improves performance on certain reasoning benchmarks Aryabumi et al. (2024); Petty et al. (2024), meta-features of data such as age and the use of toxicity filters affect performance on many QA tasks Longpre et al. (2024), and the balance of multilingual data affects performance on English and other languages Chang et al. (2023); Yue et al. (2025). These works uncover valuable insights, but they tend to focus on changing only a single aspect of the training recipe while keeping the rest fixed. Although rigorous, this is costly in compute and development time. We instead ask: can we leverage past findings from open language models to examine how training decisions jointly impact downstream performance?
[…]
2 Scaling Laws
2.1 Definition
[…]
$L(N,D)=\left(\left(\frac{N_{c}}{N}\right)^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{c}}{D}\right)^{\alpha_{D}}$ (1)
Later, Hoffmann et al. (2022a) introduced a similar law, which differed in the coefficients fitted, but was also based on a power law.
[…]
2.2 Maybe it’s Not Just Scaling?
Are parameter count and number of training tokens really all that are needed to accurately predict a model’s downstream performance? Intuitively the answer is “no” – there are a number of design decisions that go into model training, and all of them could have an effect on model performance.
Model Architecture Details
While the majority of modern language models follow the transformer architecture, there are some details that differ. For instance, the variety (Zhang and Sennrich, 2019) and position (Xiong et al., 2020) of layer normalization, and the type of positional encoding (Su et al., 2021; Press et al., 2022) make significant differences in model performance. Previous work, such as Gu and Dao (2023), has demonstrated empirically that holding all other factors equal, models that make better architecture decisions (Touvron et al., 2023a) outperform those that make worse decisions (Vaswani et al., 2017).
Data Composition
In addition, data composition and quality plays a major role in the final quality of a model. For instance, past work has demonstrated that training on some quantity of code improves performance on English reasoning tasks (Ma et al., 2023). Also, work has demonstrated that filtering for “educational” content allows for more efficient learning and higher performance on knowledge-based question answering tasks (Gunasekar et al., 2023).
Task Setting
Finally, there is an interplay of all the above factors with how model performance is measured. While previous work on scaling laws has mostly measured loss values, downstream users usually care about task performance, rather than validation loss on a pretraining dataset. Although there is often a correlation between the two for many tasks, certain tasks may be harder to predict from a model’s loss alone Bhagia et al. (2024). Moreover, certain tasks exhibit pathological scaling behaviour, such as inverse or U-shaped scaling Caballero et al. (2023); Wei et al. (2023); McKenzie et al. (2024), or simply more unpredictable performance Isik et al. (2024).
We ask: can we more effectively predict the performance of LLMs by devising a new set of “laws” that are not just reliant on scaling-based factors?
3 Building a Database of Publicly-Available Language Models
To approach our research question, we built a comprehensive database of publicly available language models. Our database encompasses models spanning a wide range of sizes, from 11M to 110B parameters, [^3] Including embedding parameters. and includes only distinct pretrained base models with decoder-only architectures. [^4] By distinct, we mean that each combination of training data and model architecture decisions should be unique. Variants of the same model trained on a deduplicated dataset are counted as separate, but not variants trained on the same data but with different curricula/initializations. In this section, we describe the criteria used for model inclusion, how we featurized the models, as well as the evaluation suite we used.
3.1 Data Collection
To ensure that our analysis was consistent, we applied the following criteria:
Pretrained-only:
Only base models that were pretrained from scratch were included. Fine-tuned variants, merged models, and models with additional post-training were excluded.
Architecture:
Only transformer-based decoder-only models were included to maintain uniformity. Mixture-of-experts (MoEs) or other architectures were excluded.
Publicly available information:
Only models with publicly available metadata, documented through configuration files or papers, were included. In particular, both the total number of parameters and total number of tokens trained on were required for inclusion. A full list of models and model families can be found in Appendix A.
3.2 Characterizing Models and Data
Figure 2: Taxonomy of pretraining data categories. We sorted data sources into this taxonomy based on model documentation.
We represent each model by the architectural choices it makes, as well as its choice of pretraining data. Formally, let $\mathcal{A}$ be the set of features related to model architecture, and $\mathcal{D}$ be the set of features related to the model’s pretraining dataset. For each task $T$ we want to approximate a model $M$’s true score $s_{T}$ with a prediction $\widehat{s_{T}}$:
$\widehat{s_{T}}(M)=f_{\theta}([\mathcal{A}_{M};\mathcal{D}_{M}]).$ (2)
[…]
In total, we document 92 open models along the dimensions of model features, high-level dataset features, and features derived from that model’s no-context generations. For the full set of features and definitions, please see Appendix B.
3.2.1 Features from Model Documentation
We first collect information about each model by reading source papers/blogs when available (see Appendix A for original citations), as well as data listed on the Hugging Face Hub Wolf et al. (2020).
Architectural Features:
These features capture design decisions that determine model structure. For example, total parameters (including embedding parameters), the number of transformer layers, the embedding and feed-forward dimensions, and details such as the type of layer normalization or attention variant used.
Data Features:
These features summarize pretraining data composition. Representative examples include total tokens trained on and the percentage breakdown of tokens sourced from various domains defined in Figure 2, as well as the proportion of English-language tokens. Our pretraining data domains were derived from common subdomains in open pretraining datasets Gao et al. (2020); Soldaini et al. (2024). We use the top level domains (web, code, books, reference, academic) as this tends to be the granularity at which data composition is described in papers.
3.2.2 Exploring Data Composition via Generation
Although many models do document some details of their data composition, relatively few release their full pretraining corpus. Further, even when data composition statistics are provided, they are often presented at a high level of granularity. As a result, many models in our study have missing values for data composition.
To address this, we further propose an approach of estimating the data a model was trained on by generating from the model with a context containing just the beginning-of-sequence (BOS) token for that model (or end-of-sequence if that model lacked a BOS token). We use temperature-based sampling with $T=1$, as this should in principle recover $P_{raw}(x_{t}\mid x_{<t})\approx P(x_{t}\mid x_{<t})$ in the limit of sampling infinitely from an LM that captures its distribution perfectly.
We call this free-generation. However, this has important caveats: we cannot actually sample infinitely, and a model does not reflect its pretraining data perfectly. Therefore, we use this as a “fingerprint” of the model’s training, but do not claim that this reflects a model’s pretraining dataset exactly. In practice, we sample between 5-10k free-generations from each model (5k for larger models). In addition to categorizing them (see Appendix E) by the domains in Figure 2, we also extract lower-level features:
Low-level data features:
We aggregate per-generation statistics that reflect data quality and linguistic structure. Examples include the words per sentence, the depth of the constituency tree for natural language text, and dependency length.
3.3 Evaluation Datasets and Metrics
Commonsense Reasoning / NLI
ANLI Nie et al. (2020) $\sim 163$k Brier Score
HellaSwag Zellers et al. (2019) $\sim 70$k Accuracy
Winogrande Sakaguchi et al. (2019) $\sim 44$k Accuracy
XNLI Conneau et al. (2018) $\sim 2.5$k Brier Score
Math / Logic
GSM8K Cobbe et al. (2021) 8 00080008\,0008 000 Accuracy
LogiQA2 Wang et al. (2020) $\sim 8$k Brier Score
MathQA Saxton et al. (2019) $\sim 37$k Brier Score
General Knowledge
ARC Challenge Clark et al. (2018) $\sim 2.6$k Accuracy
Lambada Paperno et al. (2016) $\sim 10$k Accuracy
MMLU Hendrycks et al. (2020) $\sim 2.85$k Accuracy
Other
TruthfulQA Lin et al. (2021) $817$ Accuracy
HumanEval Chen et al. (2021) $164$ Accuracy
Table 1: Overview of LM evaluation datasets with approximate sample counts, citations, and evaluation metrics. Datasets ANLI, XNLI, LogiQA2, and MathQA use Brier Score, while the others use Accuracy.
To assess how design choices affect reasoning capabilities, we evaluated model performance on a curated suite of datasets derived from the first version of the Open LLM leaderboard Myrzakhan et al. (2024), which was designed to capture diverse facets of reasoning (Table 1). [^5] Arithmetic and Minerva math (Brown et al. (2020); Hendrycks et al. (2021)) tasks were also initially included in this leaderboard, but we excluded them as we focused solely on base (not instruction tuned) models, and very few were able to achieve non-zero scores. We collect results for some models directly from the leaderboard, and for models not on the leaderboard we use the Eleuther LM eval harness (Gao et al., 2023) to conduct evaluations with exactly the same setting. In addition, if there were multiple versions of a task or sub-tasks, we evaluated all of them and averaged them to get the overall task score. For the full list of evaluation datasets and settings, see Appendix C.
For an evaluation dataset $T$ where the $i$-th sample is $y_{i}$ and model $M$, we define $s_{T}(M)$ with:
Accuracy
We use unnormalized, exact-match accuracy $s_{T,\text{acc}}=\frac{1}{|T|}\sum_{i=1}^{|T|}\mathds{1}{\{y_{i}=\hat{y_{i}}\}}$ for the majority of tasks. We use pass@1 for Humaneval, but group it with accuracy tasks for convenience.
Brier score
For tasks where smaller models struggle to achieve non-zero accuracy, we follow Schaeffer et al. (2023) in using multiclass brier score as an alternate continuous metric for multiple-choice tasks Brier (1950). For a task with $K$ classes, let $p_{ik}$ be the predicted probability for class $k$ on sample $i$. Then $s_{T,BS}=\frac{1}{|T|}\sum_{i=1}^{|T|}\sum_{j=1}^{K}(p_{ik}-\mathds{1}{\{y_{i}=k\}})$. [^6] Note that lower is better for brier score. Multiclass brier score ranges between 0-2.
Figure 3: Performance of plotted against their total parameters and tokens. The background colour represents Equation 1 fitted to the task, and the marker colours indicate true performance. Some tasks have different performance trends with scale. Within each task, individual models may also perform unexpectedly.
3.4 Heterogeneity in Task-specific Scaling
Before adding in other factors, we examine differences in scaling along $N$ and $D$ between our selected tasks. We fit a Kaplan et al. (2020) style law to each task. As seen in Figure 3, we see that different tasks may exhibit marked differences both in how well they follow scaling trends, as well as their individual scaling contours. For instance, TruthfulQA appears to exhibit U-shaped scaling, while Humaneval has more “outlier” models. A full list of $R^{2}$ values for tasks can be found in Appendix D.
4 Predictive Modeling
Next, given our database we fit a regressor to try to predict performance. In traditional scaling laws, regressors are fit based on power laws. However, we are now dealing with a larger number of features, some of which may not be captured well by simple parametric forms. Hence, we follow previous work on performance prediction (Xia et al., 2020; Ye et al., 2021) utilizing tree-based regressors based on XGBoost Chen and Guestrin (2016). [^7] We also performed preliminary experiments with LightGBM Ke et al. (2017) but it yielded very similar results in both prediction accuracy and feature importance.
For each evaluation benchmark, we train a model to predict the performance metric on that task based on architectural features $\mathcal{A}$ and data features $\mathcal{D}$. For each task setting, we perform 3-fold cross-validation due to the relatively small number of models, with a nested inner cross-validation over the training set in each fold. The inner cross-validation conducted grid search over a small set of hyperparameters, allowing the model to slightly vary per task. See Appendix G for more details.
Evaluation
To evaluate the predictors, we use Mean Absolute Error averaged across all models and folds. In other words, for a task with $N$ models evaluated, $\text{MAE}_{T}=\frac{1}{|T|}\sum_{i=1}^{N}|s_{T}(M_{i})-\widehat{s_{T}}(M_{i})|$. We compare the scaling-laws predictor as well as the all-features predictor against each other, but also against the median baseline, which simply predicts the median score of the models in the training set for each model in the test set of that fold.
Iterative Feature Selection
As the full set of features is very large, we sequentially selected features from the full set greedily based on which reduced MAE the most, averaged across 5 random seeds. Features were added until no reduction of at least $1\times 10^{-4}$ was observed. We started using only the two scaling laws features, and refer to this as the scaling-laws model, though it does not have the form of a traditional power law.888 Typically, scaling laws are used to extrapolate the performance of larger models. Because we use a decision-tree based predictor, our approach is less likely to extrapolate, a trade-off we opted to take to incorporate the array of scale-independent features we have, not all of which are numeric. Therefore, we moreso focus on interpolating performance within the size boundaries that we have (roughly 10M-100B parameters, and 50B-3T tokens). Examining results with a variety of other prediction methods is an interesting direction for future work. By then incorporating additional architectural or data features, we can then directly quantify the incremental predictive power afforded by these extra features. We refer to the model with the set of features as the all-features model. In all cases, we ran models with the same hyperparameter grid and the same random seeds and splits.
Significance Testing
Because the relative difference between baselines is small, we test both predictors across many seeds (50). We then ran paired t-tests on the overall MAE values for each seed, and corrected for multiple comparisons across tasks with the False Discovery Rate Benjamini and Hochberg (1995).
Benchmark Setting Baseline MAE Scaling Laws MAE All Features MAE p-val (corrected)
Accuracy
Arc Challenge 25-shot 13.23% 4.36% $\pm$ 0.12% 3.67% $\pm$ 0.09% $4.89\times 10^{-19}$
GSM8k 5-shot 15.65% 6.04% $\pm$ 0.21% 5.10% $\pm$ 0.23% $6.49\times 10^{-14}$
Hellaswag 10-shot 12.26% 3.93% $\pm$ 0.13% 3.18% $\pm$ 0.09% $6.66\times 10^{-20}$
Humaneval 0-shot 11.79% 8.08% $\pm$ 0.22% 6.93% $\pm$ 0.22% $1.46\times 10^{-12}$
Lambada 0-shot 16.89% 9.51% $\pm$ 0.33% 6.85% $\pm$ 0.25% $2.87\times 10^{-22}$
MMLU (0-shot) 0-shot 11.98% 4.76% $\pm$ 0.20% 4.10% $\pm$ 0.17% $6.02\times 10^{-13}$
MMLU (5-shot) 5-shot 12.25% 3.97% $\pm$ 0.18% 3.54% $\pm$ 0.14% $2.09\times 10^{-10}$
TruthfulQA 0-shot 3.72% 2.75% $\pm$ 0.08% 2.29% $\pm$ 0.06% $1.27\times 10^{-17}$
Winogrande 5-shot 10.14% 3.39% $\pm$ 0.08% 3.09% $\pm$ 0.07% $6.02\times 10^{-13}$
Brier score
XNLI 0-shot 7.22 6.68 $\pm$ 0.11% 6.30$\pm$ 0.11% $3.16\times 10^{-9}$
ANLI 0-shot 5.90 6.74 $\pm$ 0.19% 6.53$\pm$ 0.21% $3.84\times 10^{-4}$
MathQA 0-shot 7.57 2.83 $\pm$ 0.06% 2.75$\pm$ 0.07% $1.63\times 10^{-4}$
LogiQA2 0-shot 12.74 4.74 $\pm$ 0.12% 4.60$\pm$ 0.15% $1.37\times 10^{-2}$
Table 2: Comparison of MAE values (mean $\pm$ 95% CI) for Scaling Laws and All Features predictors alongside Baseline MAE. Lower MAE is bolded; * indicates significance ($p<0.05$). Brier score values are multiplied by 100 to be on a similar scale to accuracy.
5 Results
5.1 Predictor Performance
Incorporating scale-independent features consistently improves benchmark performance. We find that incorporating extra features alongside traditional scaling laws features leads to substantial improvements in prediction accuracy across multiple benchmarks, as seen in Table 2. The all-features predictor outperforms the scaling-laws-only predictor in all evaluated cases, with improvements ranging from approximately 3% (MathQA) to about 28% (Lambada) relative error reduction. Notably, the strongest improvements were observed in language modeling and common-sense reasoning tasks.
Certain tasks are more strongly dependent on non-scale features. This pattern of improvements suggests that architectural and training data features may be more informative for predicting performance on certain types of tasks more strongly linked to a particular “genre” of data. Large improvements were observed for both code generation (HumanEval, $15\%$ improvement) as well as natural-language based reasoning tasks (e.g. Lambada, $28\%$ improvement). Even tasks with narrower domains, such as mathematical reasoning (GSM8k, +16%) or knowledge-intensive evaluations (MMLU, +11–14%), see consistent, if more moderate, enhancements. The Brier score benchmarks, however, show smaller improvements (around 3–6%). This may be because the Brier score is inherently less sensitive to emergent effects in model performance, the specific choice of tasks limits the room for improvement, or a combination of both factors.
Feature Name Description [D] Total Tokens (B) Total number of tokens used during pretraining, measured in billions (log scale). [A] Total Parameters Total number of parameters in the model (log scale). [D] % Code in Pretraining Percentage of pretraining data that consists of code. [F] Question Words Ratio Ratio of question-related words generated by the model. [A] Dimension Embedding dimension. [A] Sequence length Sequence length. [A] LayerNorm Type of layer normalization used (non-parametric, parametric, rmsnorm). [A] Biases Presence of bias parameters in the model. [A] Positional Embeddings Type of positional embedding used (alibi, learned, rope). [F] % English Generated Percentage of English text generated by the model. [D] % Academic in Pretraining Percentage of pretraining data from academic sources. [D] % Reference in Pretraining Percentage of pretraining data from reference sources. [F] % Generated Weblike Text Percentage of web-like text generated by the model. [F] % Generated Reference Text Percentage of reference-like text generated by the model. [D] % Books in Pretraining Percentage of pretraining data from books. [D] % English in Pretraining Percentage of English text in the pretraining data.
Figure 4: In all tasks, the number of parameters and pretraining tokens heavily influences the predictions made by the regressor. The percentage of code in pretraining often influences predictions negatively for NLI tasks but positively for Humaneval. [D], [A] and [F] denote features derived from data, architecture, or free-generations of a model respectively.
5.2 What Features Does Task Performance Depend On?
To better understand the factors that influence task performance, we examine Shapley (1953) (SHAP) values of the predictor, which provide a local view of how individual feature values influence predictions. The results and feature descriptions for Arc Challenge, HumanEval, Winogrande, and TruthfulQA are shown in Figure 4, and results for remaining benchmarks are shown in Appendix I.
A little code goes a long way, but too much is harmful to NLI. One of the most important non-scaling features is the percentage of code data in pretraining. Higher code composition results in positive Shapley values (i.e. higher predicted performance) for Humaneval, but it negatively affects Arc Challenge, Hellaswag, Winogrande, and Lambada. In the scatterplots of Figure 5, we compare the code percentage against SHAP impact for both small and large models. We see that models trained with more than roughly 20–25% code are predicted to have large gains on tasks like Humaneval, but start to incur penalties on standard natural language benchmarks. By contrast, a moderate code proportion in the 15–25% range appears to balance these competing demands, yielding a more neutral or slightly positive effect overall.
Figure 5: SHAP impact of code percentage on Lambada (reprentative NL task) and Humaneval on our regressors.
Other domains of data can occasionally yield task-specific effects. While the percentage of code in pretraining is consistently selected as an impactful feature, with clear trade-offs, this other pretraining domains are selected less frequently. From the fine-grained features from free generations, we also observed that many recent models (particularly those trained on synthetic data such as the Phi (Gunasekar et al., 2023) and SmolLM (Allal et al., 2024)) generate a relatively large number of question words, indicating extensive training on data related to question answering. A higher percentage of reference-like or question-loaded generations resulted in better model accuracy on some tasks such as Arc Challenge and Winogrande. Additionally, models that generate more web-like data tend to do worse on TruthfulQA (Figure 4).
Non-scale architectural decisions have minor effects. Most highly influential features were data-related or architectural features related to scale (e.g., dimension). However, both the type of layer norm and the positional embedding were deemed to have a significant effect in some cases.
6 Related Work
6.1 Empirical Data Composition Results
Prior work has examined the effects of including code during pretraining Ma et al. (2023); Aryabumi et al. (2024) and ablating domains such as C4 or books from The Pile Longpre et al. (2024). Data filtering has also been shown to improve performance beyond scaling alone by removing low-quality data Sorscher et al. (2023); Goyal et al. (2024). Our results align with prior findings, indicating that code can enhance natural language reasoning at moderate proportions but degrades performance at higher percentages. We estimate an optimal code ratio of $15\text{-}25\%$, refining prior work suggesting $25\%$ Aryabumi et al. (2024), though intermediate ranges were not tested in their study. Our approach—first pooling insights from existing models—complements empirical ablations by identifying useful axes of variation for further testing.
6.2 Observational and Task-Specific Scaling Law Fitting
Previous works have examined task-specific scaling laws. In machine translation, parameter allocation between encoder and decoder affects outcomes, and incorporating machine-translated data can be detrimental Ghorbani et al. (2021). Multilingual studies reveal that language similarity doesn’t impact scaling trends; however, multitasking offers greater benefits when English is the target language Fernandes et al. (2023). Scaling laws for downstream tasks and transfer learning have been proposed, emphasizing that alignment between pretraining data and downstream tasks is crucial for performance prediction Hernandez et al. (2021); Isik et al. (2024). Data repetition has been considered, especially in data-constrained domains Muennighoff et al. (2024), with extensions to multiple data domains Goyal et al. (2024). Alternative scaling formulations address factors like sparsity Frantar et al. (2023), precision Kumar et al. (2024), and inference costs Hoffmann et al. (2022b). In contrast, some studies find stable performance across various batch sizes and learning rates DeepSeek-AI et al. (2024a).
Ruan et al. (2024) also use observations from open-source models to predict task performance, but derive their predictions of one task’s performance from performance on other tasks. We find a similar result in identifying two axes of performance– general natural language ability and coding ability but are motivated instead by tracing these capabilities back to pretraining decisions.
6.3 Pretraining Data Selection
Domain mixing has been studied in pretraining, and other works have formulated this as a regression problem Ye et al. (2024); Liu et al. (2025) or used proxy models to select domain weights in the course of training Xie et al. (2023); Albalak et al. (2023); Jiang et al. (2024b); Yu et al. (2025). In contrast, we retrospectively analyze how domain composition and training decisions influence performance across tasks, which is a complementary perspective to optimizing data weights for a single model during training.
6.4 Tracing Capabilities to Data
Specific language model capabilities have been linked to patterns in pretraining data. Performance on numerical reasoning and syntactic rule learning depends on frequency of numerical terms in the training data (Kassner et al., 2020; Wei et al., 2021). Ruis et al. (2024) found that influential data for reasoning is dispersed across numerous documents and is associated with procedural content. Similarly, Chen et al. (2024) observed that "parallel structures" are closely tied to in-context learning abilities. We currently focus on broader data domains, but our framework can be extended with more granular tasks or refined data features.
7 Conclusion and Future Work
[…]
Limitations
Our current work has several limitations that can be improved in future work. First, although we document many open models, our sample size remains limited, particularly for larger ($>$50B) parameter models. This limits our ability to draw robust conclusions about scaling behaviour in large models. Additionally, the models that we have are not evenly distributed across number of parameters, data size, and data distributions, with certain size ranges and data distributions being overrepresented. There are also likely selection effects in which models are made open-weights, as well as likely time effects in popular architectural decisions or data compositions in different time periods.
Second, our methodology also imposes some limitations. Because we do not systematically train all our own models (though we have a few of our own in Appendix A), our analyses are observational in nature. While we can observe interesting relationships between design choices and performance, making causal claims requires experimental validation. Additionally, while tree-based regressors are effective for capturing complex feature interactions, they limit our ability to extrapolate beyond the range of model sizes (in parameters and tokens) seen in our dataset.
Last, we note that the scope of our work also has limitations. Namely, we focus on base pretrained decoder-only dense transformer models, which excludes significant architectural variants such as mixture-of-experts models, non-transformer based architectures, as well as post-trained models. Additionally, we examine mostly English-language models as we do not focus on multilinguality in this work. Our feature set, while extensive, may also not capture all relevant details of model design and training, particularly optimization details as of now.
These limitations suggest directions for future work: expanding the database to include more diverse model types and language coverage, developing more targeted functional forms that allow better extrapolation while also taking as input a heterogeneous feature set, as well as conducting targeted experiments with new pretrained models to validate the impact of specific design choices.
Ethical Considerations
In this work, we focus on understanding why models may perform well on standard benchmarks, but do not focus on other important considerations such as safety or societal bias.
Furthermore, our analysis focuses on English-language models and benchmarks. This limitation reflects but may also reinforce the field’s existing bias toward English, potentially contributing to underinvestment in developing effective architectures for other languages.
Author Contribution Statement
We follow a slightly modified version of the CRediT author statement.
  • EL: conceptualization (lead), data curation (equal), methodology/software (lead), writing – original draft (lead), writing – review and editing (lead), visualization (lead), funding acquisition (supporting)
  • AB: conceptualization (supporting), methodology/software (supporting), data curation (equal), writing – review and editing (supporting)
  • LS: conceptualization (supporting), methodology/software (supporting), data curation (equal)
  • LT: conceptualization (supporting), methodology/software (supporting), data curation (equal)
  • PF: conceptualization (supporting), methodology/software (supporting), data curation (equal), writing – review and editing (supporting)
  • LM: methodology/software (supporting), data curation (equal)
  • MC: methodology/software (supporting), data curation (equal)
  • SS: methodology/software (supporting), data curation (equal)
  • CL: writing – review and editing (supporting), project administration (supporting)
  • AR: conceptualization (supporting), writing – review and editing (supporting), supervision (supporting)
  • KG: conceptualization (supporting), writing – review and editing (supporting), supervision (supporting), project administration (supporting)
  • GN: conceptualization (lead), methodology/software (supporting), writing – review and editing (supporting), supervision (lead), project administration (lead), funding acquisition (lead)
Acknowledgments
This work was supported by a fellowship from NEC Laboratories Europe and the BRIDGE CMU-AIST research project. EL was supported by the Natural Science and Engineering Research Council of Canada (NSERC), [funding reference number 578085]. AB was supported by a grant from the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2140739. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
We thank Xiang Yue for providing useful feedback and advice on this work, as well as Simran Khanuja for providing helpful edits. We also thank Brendon Boldt and David Mortensen for help with   formatting.
Appendix A List of all models
All models are listed in LABEL:tab:all_models:
Table 3: Model Parameter Counts by Organization (sorted by size)
Organization Model Name Parameters
EleutherAI Biderman et al. (2023) pythia-14m 14M
EleutherAI pythia-70m-deduped 70M
EleutherAI pythia-70m 70M
facebook Meta AI (2022a) opt-125m 125M
EleutherAI Black et al. (2021) gpt-neo-125m 125M
HuggingFaceTB Allal et al. (2024) SmolLM-135M 135M
EleutherAI pythia-160m 160M
EleutherAI pythia-160m-deduped 160M
None (this paper) llama2_220M_nl_100_code_0 220M
None (this paper) llama_220M_nl_80_code_20 220M
None (this paper) llama2_220M_nl_40_code_60 220M
None (this paper) llama2_220M_nl_20_code_80 220M
None (this paper) llama2_220M_nl_0_code_100 220M
Salesforce Nijkamp et al. (2023) codegen-350M-mono 350M
Salesforce codegen-350M-multi 350M
Salesforce codegen-350M-nl 350M
facebook opt-350m 350M
HuggingFaceTB SmolLM-360M 360M
EleutherAI pythia-410m-deduped 410M
EleutherAI pythia-410m 410M
facebook Meta AI (2022b) xglm-564M 564M
EleutherAI pythia-1b-deduped 1B
bigscience BigScience Workshop et al. (2023) bloom-1b7 1B
EleutherAI pythia-1b 1B
cerebras Cerebras Systems (2023) Cerebras-GPT-1.3B 1.3B
microsoft Li et al. (2023) phi 1.5 1.3B
EleutherAI gpt-neo-1.3B 1.3B
EleutherAI pythia-1.4b 1.4B
EleutherAI pythia-1.4b-deduped 1.4B
HuggingFaceTB SmolLM-1.7B 1.7B
Salesforce codegen-2B-mono 2B
Salesforce codegen-2B-nl 2B
Salesforce codegen-2B-multi 2B
google Gemma Team et al. (2024b) gemma-2-2b 2B
cerebras Cerebras-GPT-2.7B 2.7B
EleutherAI gpt-neo-2.7B 2.7B
NinedayWang Xu et al. (2022) PolyCoder-2.7B 2.7B
facebook opt-2.7b 2.7B
microsoft Abdin et al. (2023) phi 2 2.7B
EleutherAI pythia-2.8b 2.8B
EleutherAI pythia-2.8b-deduped 2.8B
facebook xglm-2.9B 2.9B
Qwen Qwen et al. (2025) Qwen2.5-3B 3B
cerebras Dey et al. (2023) btlm-3b-8k-base 3B
openlm-research Geng and Liu (2023) open_llama_3b_v2 3B
rinna Sawada et al. (2024) bilingual-gpt-neox-4b 4B
Dampish StellarX-4B-V0 4B
facebook xglm-4.5B 4.5B
Salesforce codegen-6B-multi 6B
EleutherAI Wang and Komatsuzaki (2021) gpt-j-6b 6B
Salesforce codegen-6B-nl 6B
Salesforce codegen-6B-mono 6B
cerebras Cerebras-GPT-6.7B 6.7B
facebook opt-6.7b 6.7B
EleutherAI pythia-6.9b-deduped 6.9B
EleutherAI pythia-6.9b 6.9B
Qwen Bai et al. (2023) Qwen-7B 7B
aisingapore Lowphansirikul et al. (2021) sea-lion-7b 7B
bigscience bloom-7b1 7B
google Gemma Team et al. (2024a) gemma-7b 7B
mosaicml MosaicML NLP Team (2023) mpt-7b 7B
openlm-research open_llama_7b 7B
tiiuae Institute (2023) falcon-7b 7B
allenai for AI (2024) OLMo-7B-hf 7B
huggyllama Touvron et al. (2023a) llama-7b 7B
LLM360 Liu et al. (2023) Amber 7B
LLM360 CrystalCoder 7B
facebook xglm-7.5B 7.5B
meta-llama AI (2023) Meta-Llama-3-8B 8B
google gemma-2-9b 9B
01-ai 01. AI et al. (2025) Yi-9B 9B
EleutherAI pythia-12b 12B
EleutherAI pythia-12b-deduped 12B
cerebras Cerebras-GPT-13B 13B
meta-llama Touvron et al. (2023b) Llama-2-13b-hf 13B
Qwen Qwen1.5-14B 14B
Qwen Qwen2.5-14B 14B
Salesforce codegen-16B-nl 16B
Salesforce codegen-16B-mono 16B
EleutherAI gpt-neox-20b 20B
mosaicml mpt-30b 30B
Qwen Qwen2.5-32B 32B
Qwen Qwen1.5-32B 32B
AbacusResearch Jallabi-34B 34B
01-ai Yi-34B 34B
01-ai Yi-34B-200K 34B
meta-llama Llama-2-70b-hf 70B
meta-llama AI (2023) Meta-Llama-3.1-70B 70B
meta-llama Meta-Llama-3-70B 70B
Qwen Qwen et al. (2025) Qwen2-72B 72B
Qwen Qwen2.5-72B 72B
Qwen Qwen1.5-110B 110B
Appendix B List of all architectural and data features
B.1 Architectural Features
Note that features in this section are collected from official documentation (e.g. huggingface model/data cards or original papers).
  • Total parameters - the total number of parameters (embedding included) in the model. Note that we only include decoder-only dense models.
  • Dimension - the embedding dimension.
  • Num heads - the number of attention heads.
  • MLP ratio - the ratio of FFN dimensionembedding $\frac{\text{FFN dimension}}{\text{embedding dimension}}$.
  • Positional Embeddings - the type of positional embedding. This is either non-parametric (sinusoidal or fixed embeddings), learned (just learned as a vector per position), rope (rope embeddings), or alibi (technically not an embedding, but included here due to its functional purpose)
  • LayerNorm - the type of layernorm applied. This is either non-parametric (just an arithmetic based normalization), parametric (similar, but with some learnable parameters such as scaling/biases), and RMSNorm (a simplified version of parametric)
  • Attention variant - The broad type of attention used. This is either full (vanilla attention), local (each token position only attends to positions around it), mqa (multi-query attention), or gqa (grouped-query attention)
  • Biases - whether or not bias terms are present in parts of the model. Either none (no biases), attn only (only in attention layers), ln only (only in layer norm)
  • Block type - whether or not the transformer blocks are computed in parallel at all. Sequential indicates not, while parallel indicates some parallelism in attention or FFN layers.
  • Activation - the activation function used. Either relu, gelu/gelu variations, silu, or swiglu.
  • Sequence length - the sequence length.
  • Batch instances - the batch size used during pretraining.
B.2 Data Features
Note that features in this section are collected from official documentation (e.g. huggingface model/data cards or original papers).
  • Total tokens (B) - total number of tokens used during pretraining, measured in billions (converted to log scale)
  • % Web in Pretraining - Percentage of pretraining data from general web sources.
  • % Code in Pretraining - Percentage of pretraining data that consists of code.
  • % Books in Pretraining - Percentage of pretraining data from books.
  • % Reference in Pretraining - Percentage of pretraining data from reference sources.
  • % Academic in Pretraining - Percentage of pretraining data from academic sources.
  • % English in Pretraining - Percentage of English text in the pretraining data.
B.3 Freegen-derived Features
These features are derived from model generations. For each model, 5–10k generations are extracted and the following metrics are aggregated (by mean and standard deviation). However, bigram entropy, the educational classifier score, and domain classifications are exceptions, as they are computed once across all generations.
We use Stanza Qi et al. (2020) to generate the parse-based features after classifying generations by language. We only include languages that are supported by stanza in the final set of generations that the parse features are based on.
B.3.1 Generation Length & Basic Statistics
  • Mean Character Length – Average number of characters per generation (capped at 2048).
  • Mean Tokens Generated – Average number of tokens per generation.
  • Mean Sentences – Average number of sentences per generation.
  • Mean Words – Average number of words per generation.
  • Mean Words per Sentence – Average number of words per sentence.
B.3.2 Constituency Parse Features
  • Mean Depth of Deepest Parse Tree – Average maximum constituency tree depth per generation.
  • Mean Depth of Parse Trees – Average constituency tree depth across all sentences/phrases.
  • Mean Word Depth – Average depth of words within constituency trees.
  • Mean Word Depth Variation – Average standard deviation of word depths across sentences/phrases.
B.3.3 Dependency Parse Features
  • Mean 90th-Percentile Dependency Head Distances – For each generation, compute the 90th-percentile of the linear distances between words and their dependency head, then average these values.
  • Mean Maximum Dependency Head Distances – Average maximum distance from any word to its dependency head per generation.
  • Mean Median Dependency Head Distances – Average median dependency-head distance per generation.
  • Mean Maximum Dependency Root Distances – Average maximum distance from any word to the sentence root per generation.
  • Mean Mean Dependency Root Distances – Average of the mean distances from words to the sentence root per generation.
  • Mean Median Dependency Root Distances – Average of the median distances from words to the sentence root per generation.
B.3.4 Domain Classification Features
  • % Generated Academic-like Text – Percentage of generations classified as academic-like.
  • % Generated Books-like Text – Percentage of generations classified as books-like.
  • % Generated Code-like Text – Percentage of generations classified as code-like.
  • % Generated Reference-like Text – Percentage of generations classified as reference-like.
  • % Generated Specialized Text – Percentage of generations classified as specialized (e.g., music scores, chess PGNs, biomedical data).
  • % Generated Web-like Text – Percentage of generations classified as web-like.
B.3.5 Classifier and Language Metrics
  • Mean Educational Classifier Score – Average score assigned by the educational classifier.
  • % Generated English Text – Average percentage of text generated in English.
B.3.6 Lexical Diversity and Entropy Metrics
  • Mean Bigram Entropy – Average entropy computed on bigrams across generations.
  • Type-Token Ratio – Average ratio of unique tokens to total tokens.
  • Unique Tokens – Average number of unique tokens per generation.
B.3.7 Lexical and Stylistic Features
  • Content-Function Ratio – Ratio of content words (nouns, verbs, adjectives, adverbs) to function words.
  • Question Words Ratio – Ratio of question-related words (e.g. how, what, why, when, where, who, which, whose) per 100k words.
  • Imperative Words Ratio – Ratio of imperative words (e.g. do, make, consider, take, use, ensure, check, build, apply, run, create, find, go, try, turn, start, stop, put, keep, leave, get, move) per 100k words.
  • Conjunctions Ratio – Ratio of conjunction words (e.g. and, but, or, so, because, although, however, therefore, yet) per 100k words.
  • Instruction Words Ratio – Ratio of instruction-oriented phrases (e.g. “Question:”, “Answer:”, “Instruction:”, “User:”, “Assistant:”, “Q:”, “A:”) per 100k words.
  • Numbers Ratio – Ratio of numerical tokens in the generated text.
Appendix C List of all evaluations and settings
Although we ideally would evaluate the full cross-product of models and tasks, we found that due to some models being incompatible with LM Evaluation Harness and compute constraints we could not evaluate all 92 models on every dataset. We list in Table 4 the number of evaluations we currently have for each benchmark and will continue to fill out evaluations in the database.
Task Name # Models Evaluated
Commonsense Reasoning / NLI
ANLI 82
HellaSwag 92
Winogrande 92
XNLI 82
Math / Logic
GSM8K 92
LogiQA2 82
MathQA 82
General Knowledge
ARC Challenge 92
Lambada 92
MMLU 92
Other
TruthfulQA 92
HumanEval 91
Table 4: Number of models evaluated for each benchmark task. Note that some models encountered technical errors when being loaded or in lm-eval-harness. The number of models will continue to be updated.
Appendix D Task Deviations from Kaplan-style Scaling Laws
In Table 5, we document the $R^{2}$ value for a fitted power law on the performance of each model.
Benchmark $R^{2}$
gsm8k 0.85
arc challenge 0.82
hellaswag 0.80
winogrande 0.80
mmlu 5-shot 0.80
mmlu 0-shot 0.74
mathqa 0.70
ANLI 0.61
humaneval 0.61
lambada 0.55
LogiQA2 0.50
XNLI 0.41
truthfulqa 0.29
Table 5: Overview of $R^{2}$ values by benchmark.
Appendix E Free-generation Domain Classification
We classify model generations into top-level domains with GPT-4o-mini. We found that this multi-stage prompt 1, 2 had reasonable precision on a sample of Dolma by domain Soldaini et al. (2024), so use it to classify free-generations.
Listing 1: Multistage classification prompt.
<PROMPT 1>. [SYSTEM] You are a system tasked with classifying documents. First, determine if this document is relatively coherent. These documents are generated by language models, so they may not make sense. Classify a document as incoherent ONLY if it shows extreme repetition, code mixes in a way that does not make sense (such as different languages referencing entirely different subjects), or if it is mostly gibberish. Dont worry about logic errors or factual inconsistencies. If multiple documents are mixed into one, classify it as incoherent. Respond ONLY with "incoherent" if the document is incoherent, otherwise respond with "not_incoherent"
[…]
If not incoherent...
<PROMPT 2>. [SYSTEM] Determine if this document contains programming code. Look for:
1. Programming language keywords (def, class, import, etc)
2. Code blocks (marked with backticks, indentation patterns)
3. Stack Overflow-style Q&A about programming
4. File extensions (.py, .js, etc)
5. Documentation about code/config files
Respond ONLY with:
- "code" if ANY of these are present
- "not_code" otherwise
[…]
If not code...
<PROMPT 3> [SYSTEM] For documents WITHOUT programming code, determine if this is web content. Web content includes news articles, social media and online forums, blog posts, shopping websites, and other general websites. This includes a wide variety of content, and anything that looks like it may be a web article at all should be included. Look for:
1. URLs or hyperlinks
2. Social media formatting (@mentions, #hashtags)
3. "Click here" or UI elements
4. Comment threads or forum posts
5. Shopping/e-commerce language
6. Bylines or author names
7. Descriptions of products or product features
Respond ONLY with:
- "web" if ANY of these are present
- "not_web" otherwise
[…]
If not web...
<PROMPT 4> [SYSTEM] For documents WITHOUT programming code, determine if this is academic or patent-related content. Academic content consists of research papers and snippets of research in both sciences and humanities, as well as patent applications. Student essays or assignments should also be included in this category. Look for:
1. Citations or references
2. Latex formatting such as equations or tables
3. Formal academic language, not aimed at educating a general audience
4. Technical jargon or domain-specific terminology
5. Patent numbers or legal language (but not court documents, only patents)
Do NOT classify as academic if the document:
- Only uses occasional technical terms
- Is a popular science article or description of a scientific study, rather than the study itself
- Is educational but aimed at a general audience
Respond ONLY with:
- "academic" if ANY of these are present
- "not_academic" otherwise
[…]
Listing 2: Multistage classification prompt (contd).
If not academic...
<PROMPT 5> [SYSTEM] For documents WITHOUT programming code, determine if this is a book, reference material (including media content), or a specific dataset. Books include literary works, fiction, and narrative nonfiction. Reference material includes wikipedia, dictionaries, textbooks and textbook like content, and encyclopedias. Please note that reference should also include instruction or human preference datasets for language model training. Media content includes podcasts, subtitles, and other media-related text. Specific datasets are unique and not covered by the other categories, such as biomedical datasets or molecules, chess PGNs or specific data formats not covered by any other category. Look for:
For the books category:
1. Chapter headings or book titles
2. Fictional character names or dialogue
3. For literary nonfiction, look for a more narrative and less didactic tone
4. Extended narrative prose or dialogue
Do NOT classify as books if the document:
- Only has a single dialogue snippet
- Could be a web article
- Is primarily informational or educational (use reference instead)
For the reference category:
1. Definitions or explanations of terms
2. Encyclopedic formatting
3. Textbook-like language
4. Explanations or examples meant to educate a reader
5. Chat formatting like User:/Assistant:’ or similar tokens
6. Court documents or legal language (NOT patents)
7. Wikipedia headers such as references or external links
For the media category (should be classified as reference):
1. Audio or video timestamps
2. Subtitles or captions
For the specific datasets category:
1. Unique names or identifiers
2. Dataset-specific formatting
3. Data or metadata descriptions
If this seems to be a web document (social media, news, blogs, forums, shopping), you can also back off to the web category.
Respond ONLY with:
- "books" if the document is a book
- "reference" if the document is reference material
- "specific_datasets" if the document is a specific dataset
- "web" if the document is web content
- "unknown" if none of these are present
[…]
Appendix F Free-generation Examples
Llama-2-70B, web-like: Gwen Stefani Couldn’t Recognize Her First Song at an Event with Blake Shelton What’s the first song you ever wrote and or recorded? And how would you feel if you never heard it again until you were Facetiming with hundreds of thousands of people while attending a big country music event? That’s exactly what happened to Gwen Stefani, when she was honored at the Country Radio Seminar event this week. In Nashville, . One can only imagine that when you’ve had as crazy a career as she’s had, it may be a chore to actually remember all of the milestones over the years, but just imagine also doing it on the spot, for everyone to see. Well, as it turns out, the doubly talented Dolly Parton was set to introduce Gwen as “an artist who broke barriers” when she cited the lyrics from one of Stefani’s first-ever “country” songs…“year 3,000.” The diplomatic “The Voice” coach Blake Shelton interjected at the video she was showing at that moment, asking “was there other work done ….
phi-2, books-like: The sun glistened brightly in the evening sky, casting warm rays across the quaint little kitchen where a pair of diligent parents were preparing a scrumptious meal. The air was filled with anticipation as they embarked on a culinary adventure, their hearts filled with love and joy.
With their sleeves rolled up, the parents deftly wielded an assortment of sharp knives, each one an extension of their skill and precision. Like artists with their brushes, they delicately maneuvered through the ingredients, dancing with knives as their partners. The blades moved with a swiftness that defied the physics of everyday movements, gliding through the air with grace and purpose.
Qwen2-72B, reference (instruct)-like: Children: Family ticket (at most 2 adults and 2 children) \$23.00 (1) Mr Black wants to see the film with his wife and two sons (they’re both over 9), how much should he pay?____ A. \$23. B. \$46. C. \$48. D. \$49. (2)Mr Brown wants to see the film with his 5-year-old twin sons, how much should he pay____. A. \$18. B. \$23. C. \$28.5 D. \$32 (3)Lisa can’t see snow in this film.____ A. True B. False (4). Jane is a maths teacher and she has two daughters (she’s 12 and her sister is 8). They want to see the film. How much should she pay?____ A. \$28.5 B. \$27 C. \$33 D. \$26.5 (5) Mike is sixteen. Today is Saturday and he wants to see the film with his 40-year-old parents. They will pay____ for the tickets. A. \$22.
OPT-6.7B, web-like: I don’t think this is a question. It’s a statement. I say get a dog, the world is better with dogs.
I said to myself the exact same thing. I don’t know why you were down-voted.
because reddit is a bunch of people who hate dogs.
Codegen-16B-mono, code-like:
1from typing import Optional
2
3from fastapi import FastAPI
4from pydantic import BaseModel
5
6app = FastAPI()
7
8
9class Item(BaseModel):
10 name: str
11 price: float
12 is_offer: Optional[bool] = None
13
14
15@app.get("/")
16def read_root():
17 return {"Hello": "WorldfromFastAPI"}
18
19
20@app.get("/items/{item_id}")
21def read_item(item_id: int, q: Optional[str] = None):
22 return {"item_id": item_id, "q": q}
23
24
25@app.put("/items/{item_id}")
26def update_item(item_id: int, item: Item):
27 return {"item_name": item.name, "item_id": item_id}
Pythia-410M, academic-like (patent): This invention relates to new and improved methods and apparatus for removing solid waste material from the waste stream of a power station where the solid waste material is intended for disposal after completion of the power plant.
Various attempts have previously been made to remove solid waste from the waste stream of plants. This is true, for example, to the discharge of sludge which is generally collected into a sludge tank and washed out of the plant in a washing tank which is generally connected to an open drain outlet of the plant. This prior art is discussed by U.S. Pat. No. 3,623,579 which issued to G. R. Clark and described a method for treating the waste stream to remove solid waste by flocculating and flocculating and agitating the solids in a tank to break bonds between the solid particles.
Furthermore, an apparatus was described by U.S. Pat. No. 4,016,823 which describes a method in which liquid sewage is removed from the waste stream and from the sewage treatment plant where the solid waste being removed is to be treated to produce ammonia-purified water for use in bathing baths or soaps and where the sewage from the wastewater treatment plant is removed to the sewage processing plant where this sewage is mixed with water or treated as a fertilizer.
Appendix G XGBoost Settings
For the inner grid search, the maximum depth of trees was in $[2,3,5]$, while the learning rate was in $[0.01,0.1,0.3]$ and the number of trees was in $[50,100]$.
Appendix H Selected Features by Task
In Table 6, we show the selected features per benchmark.
Table 6: Greedily-selected features per benchmark.
Benchmark Selected Features
arc challenge (25-shot) total params, pretraining summary total tokens billions, question words ratio, layer norm type, dimension, pretraining summary percentage code
gsm8k (5-shot) total params, pretraining summary total tokens billions, pretraining summary percentage reference, edu classifier std, pretraining summary percentage books
hellaswag (10-shot) total params, pretraining summary total tokens billions, pretraining summary percentage code, pretraining summary percentage reference, positional embeddings, pretraining summary percentage academic
mmlu 0-shot (0-shot) total params, pretraining summary total tokens billions, layer norm type, activation, pretraining summary percentage code
truthfulqa (0-shot) total params, pretraining summary total tokens billions, domain web pct mean, dep parse dep root dist max mean, pretraining summary percentage english, entropy mean, layer norm type
winogrande (5-shot) total params, pretraining summary total tokens billions, question words ratio, layer norm type, pct english mean, positional embeddings, pretraining summary percentage books, pretraining summary percentage code, block type
anli (0-shot) total params, pretraining summary total tokens billions, pretraining summary percentage code, pretraining summary percentage web, pretraining summary percentage books, positional embeddings
logiqa2 (0-shot) total params, pretraining summary total tokens billions, pretraining summary percentage web, domain reference pct mean, dep parse dep root dist mean std, dep parse dep root dist median std
mathqa (5-shot) total params, pretraining summary total tokens billions, pretraining summary percentage books, num heads
xnli (0-shot) total params, pretraining summary total tokens billions, pretraining summary percentage web
lambada (0-shot) total params, pretraining summary total tokens billions, pretraining summary percentage code, block type
mmlu 5-shot (5-shot) total params, pretraining summary total tokens billions, sequence length, biases, num heads, dimension, edu classifier mean, pretraining summary percentage academic
gsm8k (5-shot) total params, pretraining summary total tokens billions, instructions words ratio, pretraining summary percentage academic, sequence length, mlp ratio
humaneval (0-shot) total params, pretraining summary total tokens billions, pretraining summary percentage code, layer norm type, pretraining summary percentage english, biases
Appendix I SHAP Plots for remaining benchmarks
SHAP plots for the remaining benchmarks can be found in Figure 6Figure 14. Please note that lower scores are better for Brier score tasks (ANLI, XNLI, MathQA, LogiQA2)
Figure 6: SHAP values for GSM8k
Figure 7: SHAP values for Lambada
Figure 8: SHAP values for Hellaswag
Figure 9: SHAP values for MMLU 0-shot
Figure 10: SHAP values for MMLU 5-shot
Figure 11: SHAP values for ANLI
Figure 12: SHAP values for XNLI
Figure 13: SHAP values for MathQA
Figure 14: SHAP values for LogiQA2